zackariahharriett:

Jamie is 5 years older than Ella. Jamie's age is 11 years less than three times Ella's age. The system below models the relationship between Jamie's age (j) and Ella's age (e): j = e + 5 j = 3e − 11 Which of the following methods is correct to find Jamie's and Ella's age? (4 points) Solve j + 5 = 3j − 11 to find the value of j. Write the points where the graphs of the equations intersect the x-axis. Write the points where the graphs of the equations intersect the y-axis. Solve e + 5 = 3e − 11 to find the value of e.

draebrewer:

i know the answer

CripQUEZZ:

I’d say A D you need to jus solve it out like this we will use Ella as an example Ok so it’s set up like E + 5 = 3e - 11 First get your variable on one side So E - E = 3e - E So the first E cancels out and your left with 5 = 2e - 11 Now you wanna cancel out the 11 and get the variable by it’s self So do the opposite of what 11 is doing and add it to both sides 5 + 11 = -11 +11 So the -11 cancels out and now it’s 5 + 11 = 2e Then simplify the addition 5 + 11 = 16 So now it’s 16 = 2e Now do the opposite of what the 2 is doing to the E and divide it from both sides to cancel it out 16/2 = 2e/2 Then the 2e/2 is canceled out and left with the variable by its self 16/2 = e Then simplify 16/2 = 8 So that means E = 8

CripQUEZZ:

So now use this for the other one

zackariahharriett:

Solve e + 5 = 3e − 11 to find the value of e.

zackariahharriett:

and i picked that off of someone cant say

CripQUEZZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @zackariahharriett Solve e + 5 = 3e − 11 to find the value of e. \(\color{#0cbb34}{\text{End of Quote}}\) U just summarized exactly what I said

zackariahharriett:

actually draebrewer was the first to give e the answer in my message

CripQUEZZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @zackariahharriett actually draebrewer was the first to give e the answer in my message \(\color{#0cbb34}{\text{End of Quote}}\) So he gave the direct answer

zackariahharriett:

yea

zackariahharriett:

and explaind it in my message